OpenStudy (anonymous):

Solve for x. thx u. (4x)^lg5 = (5x)^lg7

OpenStudy (anonymous):

is this \[(4x)^{\ln(5)}=(5x)^{\ln(7)}\]

OpenStudy (anonymous):

lg5 = log base10 5

OpenStudy (anonymous):

\[(4x)^{\log(5)}=(5x)^{\log(7)}\] \[4^{\log(5)}x^{\log(5)}=5^{\log(7)}x^{\log(7)}\]

OpenStudy (anonymous):

only solution will be x = 0